### v4

# A Data Perspective on Enhanced Identity Preservation for Diffusion Personalization

Abstract
Large text-to-image models have revolutionized the ability to generate imagery using natural language. However, particularly unique or personal visual concepts, such as pets and furniture, will not be captured by the original model. This has led to interest in how to personalize a text-to-image model. Despite significant progress, this task remains a formidable challenge, particularly in preserving the subject’s identity. Most researchers attempt to address this issue by modifying model architectures. These methods are capable of keeping the subject structure and color but fail to preserve identity details. Towards this issue, our approach takes a data-centric perspective. We introduce a novel regularization dataset generation strategy on both the text and image level. This strategy enables the model to preserve fine details of the desired subjects, such as text and logos. Our method is architecture-agnostic and can be flexibly applied on various text-to-image models. We show on established benchmarks that our data-centric approach forms the new state of the art in terms of identity preservation and text alignment.
1 Introduction
Figure 1: Demonstration of our method applied to both SDXL and SD, compared with several prior works. Note that our results preserves very fine details of the subject without overfitting to the training images and losing text alignment.
Recent advancements in text-to-image models have greatly improved the creation of diverse and high-quality images based on text input [24, 27, 20]. However, these models still face challenges in accurately generating images of specific subjects, such as unique backpack designs or individual dog breeds. Text, being a vague representation, cannot fully capture the intricate details of real-world subjects. This limits the models’ capability in generating novel images of the exact subjects.
To tackle this issue, previous research has explored fine-tuning large-scale diffusion models, such as DreamBooth [25] and TextInversion [7], to associate an identifier token with the desired subject. This not only preserves subject identity but also exhibits impressive generalization across various text inputs [25, 7]. Nevertheless, these methods still face challenges in reproducing fine details of the subject, as shown in Figure 1.
We argue the reason that the previous methods fail to preserve identity details, e.g. text and logos, as shown in Figure 1, Figure 14, 15, 16 in Appendix F, is because they cannot capture enough training information [12, 13, 32, 21, 9, 33, 30].
A naive way to capture more training information is to fine-tune the model with more iterations. However, a longer training period can lead to overfitting. As a result, the model ignores text inputs and can only generate images that are visually similar to training images. To address the overfitting issue, DreamBooth [25] proposed a regularization dataset where images are generated by the text input “a [class noun]” (e.g., “a backpack”). The regularization dataset is used to fine-tune the model, along with the training images, which are prompted as “a [identifier_token] [class noun]” (e.g., “a sks backpack”). However, experiment results show that the overfitting issues still exist (Figure 5). As a result, their method has to stop the fine-tuning early and cannot preserve the fine details.
We propose a new regularization dataset generation strategy to solve the over-fitting issue. We argue that the root cause of ineffectiveness of DreamBooth lies in their over-simplified text input. The simplistic text input “a [identifier_token] [class noun]” and “a [class noun]” cannot fully describe the image information. Thus, the identifier token of the desired subject fails to distinguish between the foreground and background, and associate with the entire image.
We found that introducing the background text for training examples, “a [identifier_token] [class noun] [background]”, and generating regularization dataset with the same background text, effectively associate the identifier token with the desired subject. This approach does not associate the identifier with the entire image, and as a result, it prevents the model from overfitting. The text of the regularization dataset are generated in the format “a [shape] [color] [texture] [class noun] [background]” by Large Language Models (LLMs), and the corresponding images are generated from pre-trained diffusion models.
Our new regularization dataset generation strategy yields a more stable fine-tuning process that can continue longer without overfitting and significantly enhances both identity preservation and text alignment. As demonstrated in Figure 1, our approach excels in preserving subject-specific details while maintaining robust generalization across diverse text contexts.
This paper validates the effectiveness of our approach through extensive experimentation on the DreamBench dataset [25], which contains a broad range of subjects, including complex items like drink cans and backpacks, as well as living animals with diverse poses and movements. Our experiments encompass a range of tasks, including subject re-contextualization, attribute modification, accessorization, and style transfer. Our results confirm our method’s ability to proficiently preserve subject identity, including intricate details like logos, while demonstrating robust generalization across diverse text inputs. Additionally, we highlight limitations in the original evaluation metric CLIP-T score used in DreamBooth [25] and propose an enhanced version of CLIP-T tailored to offer more precise and informative indications of text-image alignment.
We summarize our contributions in the following aspects:
1. A data-oriented approach to tackle overfitting issues during diffusion model fine-tuning.
2. A modification to the CLIP-T score to enhance its capacity for indicating text alignment.
3. A systematic exploration of diverse methods for generating a regularization dataset, with the intention of providing valuable insights to the broader research community.
2 Related Work
Diffusion Based Text-to-Image Models have recently demonstrated remarkable progress, primarily through the utilization of diffusion models [11]. These achievements are particularly pronounced when employing large models trained on extensive datasets [24, 20, 27, 3, 18, 23]. Our method extends the pre-trained diffusion models to incorporate personalized concepts. We apply this method to well-established models, such as StableDiffusion [24] and StableDiffusionXL [20]. Note that our method is fundamentally rooted in data, and is potentially applicable to a wide range of different diffusion model architectures.
Text-to-Image Personalization aims to imbue pre-trained diffusion models with the ability to produce novel images of specific subjects [25, 7]. However, previous methods grappled with a delicate balancing act. They oscillate between underfitting the intended subject and overfitting to the training dataset and losing text alignment. [10] attempt to address the problem of underfitting the subject by injecting a reference image feature map into attention modules to preserve identity. [5], on the other hand, embark on the task of disentangling subjects from backgrounds by employing a CLIP image encoder for encoding backgrounds [22]. To combat the issue of overfitting, researchers have explored diverse strategies, mostly by fine-tuning different decomposition parts of the model [12, 13, 32, 21, 9, 33, 30]. To sidestep time-consuming fine-tuning the diffusion model, several encoder-based methods have emerged, involving the training of an additional encoder for all subjects [29, 8, 26, 2, 14]. These encoder-based methods often yield increased diversity in generated content, albeit with some trade-offs in identity preservation compared to fine-tuning-based approaches. Our approach focuses on maximizing quality at the cost of training time. It improves both identity preservation and text alignment relative to prior work. We leave the topic of accelerated training as an avenue for future exploration.
Involving More Data is our core idea. Previous work [34] has shown that a better prompt can significantly improve the generated image quality. The most similar approach to ours is proposed in DreamBooth [25]. They propose to generate a regularization dataset using prompt “a [class noun]”, e.g., “a backpack”, to alleviate overfitting. However, our experiments reveal that this generated dataset lacks the strength to effectively counter overfitting. [13] extract the real text-image pairs that similar to the training examples from existing dataset, but they still need to restrict the number of training iterations to mitigate overfitting risks. [6] speed up customization with a feed forward network trained on many datasets of custom objects, but they prioritize speed over quality improvements. [17] propose to generate masks, bounding boxes, and their corresponding tags to train a better adapter and thus a better composition method. Our work is primarily centered on enhancing generation quality, with the aspect of speed left for future exploration.
3 Method
Our main goal is to fine-tune a pre-trained text-to-image diffusion model using a limited set of subject-specific training examples, typically around 4 to 6 images. This fine-tuning process is aimed at enabling the diffusion model to generate new images of the provided specific subject.
Figure 2: Overview. In contrast to prior approaches, we introduce specific prompts to the training examples and create a regularization dataset with a wider range of images guided by these prompts. To further boost diversity of regularization dataset, we generate additional prompts and images using the same prompt formats.
Recap of DreamBooth. Instead of linking the new subject with a noun word [7], DreamBooth [25] proposed to linking the new subject with an adjective word, by inserting a special identifier token before the subject class noun. For example, “a backpack” becomes “a [identifier_token] backpack” or “a sks backpack” in DreamBooth [25]. Furthermore, to prevent overfitting, they proposed to create a dataset of text-image pairs by employing prompts in the format “a [class noun]” (e.g., “a backpack”). This dataset serves as a counterbalance to the training examples. During training, both the training examples and the regularization examples are fed into the diffusion model. However, we discovered that the divergence of the over-simplified text and the image still leads to overfitting before fine details are fully learned.
Our Approach. In contrast to previous methods using a simplistic prompt for the real training examples, we first introduced individual and more concrete prompts, such as “a backpack on a rock”. Surprisingly, this adjustment did not yield significant improvements, and the model still exhibited rapid overfitting. It extends the fine-tuning time only when combined with a strong regularization dataset using prompts that describe foreground and background in more details. Inspired by this observation, our focus shifted towards automatically constructing an appropriate regularization dataset to address the issue of overfitting.
The overview of our method is presented in Figure 2. Importantly, our approach retains the model architecture while concentrating on the development of a novel regularization dataset to combat overfitting effectively. This means that our method can be combined with different pre-trained diffusion models.
Enhancing Training with Prompts. The inclusion of prompts in training instances plays a pivotal role in our approach. While employing a uniform prompt format such as “a [identifier_token] [class noun]” for all training examples may seem straightforward, it inadvertently increases the risk of associating the token with both foreground and background, causing overfitting to the training data. To address this challenge, we developed this idea further by incorporating background descriptions into each image, as depicted in the upper-right corner of Figure 2. Additionally, we replaced the provided class name with more specific ones, which will be discussed further in Appendix B. It’s important to note that merely adding the background prompt alone does not yield the desired benefits. Instead, our contribution is on constructing a dedicated regularization dataset, which proves to be essential in addressing the issue of overfitting, as elaborated in subsequent sections.
Generating regularization examples against training prompts. We present a novel approach for generating a regularization dataset, which overcomes the ineffectiveness of the original example prompt. For instance, let’s consider the initial prompt “a backpack on a rock”. We enhance this prompt by introducing structural components in the format of “a [shape] [color] [texture] [class noun] [background]”, where [shape], [color], and [texture] represent randomly selected adjectives drawn from a pool of 100 options for each attribute. These adjectives are then combined to create prompts, such as “a contoured orchid woven backpack on a rock” and “a semicircular turquoise canvas backpack hanging on a tree branch”. The created prompts are used for image generation, as demonstrated in the lower-left corner of Figure 2. The pools of the adjectives are generated automatically by Large Language Models [19]. For more detailed information on this process, please refer to Appendix A.
Amplifying Diversity with Structured Prompts. To increase the diversity of our regularization dataset, we randomly generate 500 additional background phrases and 100 style phrases. We then combine these phrases in a random manner, resulting in structured prompts in the format “a [style] [shape] [color] [texture] [class noun] [background]”. For instance, this approach generates prompts like “a photo of a cone-shaped eggshell scaly backpack amidst the aromas of a lively outdoor market” and “a children’s storybook illustration of a trapezoidal coral embossed backpack against the canvas of a city skyline”, as showcased in the lower-right corner of Figure 2. This strategy substantially enhances image diversity compared to the simplistic use of the prompt “a [class noun]”.
Adaption for Living Entities. For living entities, we use descriptors like body, skin/fur, and emotion instead of shape, color, and texture. Additionally, we introduce motion into the backgrounds, transitioning from static scenes like “in an urban city” to dynamic contexts such as “walking in an urban city”. For a detailed explanation of this adaptation process, please refer to Appendix A.
Dropout. In our strategy to enrich the dataset, we include three descriptive words for each object, which increases diversity. However, this approach can lead to a potential issue: overfitting to sentence structure. The model might learn that one adjective word corresponds to a specific subject, while three adjective words signify a different subject. To mitigate this, we introduce randomness by randomly excluding some words during training. For example, we may have a prompt like “a [shape] [color] backpack on a rock” excluding a texture descriptor. This dropout technique adds an element of unpredictability, strengthening the model’s adaptability and reducing its reliance on fixed sentence formats.
Cropping. [32] point out that the model is prone to overfit to the image layout when attempting to learn personalized concepts from a limited set of examples. Similarly, [13] observe that employing random cropping enhances convergence speed and yields improved results. To enhance our model’s performance, we incorporate random cropping with a variable ratio ranging from 0.75 to 1. It is worth noting that for SDXL, we incorporate cropping coordinates as described by [20], where we increase the original image size to achieve a cropped image size of 1024x1024 pixels.
Implementation Details We create 2000 regularization images for each subject, distributed as follows: 20% photorealistic images have different subjects in same background; 60% photorealistic images have different subjects in different background; and 20% differently styled images have different subjects in different background. Additional details regarding the optimal iteration range can be found in Appendix E.
4 Experiments
We quantitatively evaluate on the DreamBench dataset [25], which contains 30 subjects ranging from backpacks and stuffed animals to dogs and cats. The dataset encompasses 25 distinct testing prompts for inanimate objects and 25 for living entities. Following the original paper, we generate four images per prompt, which yields a total of 3000 images for evaluation. Importantly, the testing prompts used in our experimental results are included in our regularization dataset.
Our baseline model is DreamBooth (SD backbone) [25]. For a fair comparison, we take the following steps: 1) Employ SDXL to generate 2000 high-quality regularization images; 2) Use “olis” as the identifier word; 3) Set the learning rate to 2e-6 for 1500 iterations; and 4) Apply 200 steps of DDIM for inference.
4.1 Result Comparison
The qualitative comparison is depicted in Figure 1 and Figure 10-16 in Appendix F. We use “new” to represent the “ [identifier_token]”. Our approach outperforms alternative methods in preserving intricate subject-specific details, such as logos on drink cans, unique icons on backpacks, and fine hair textures on toy monsters.
Methods Backbone DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T (vague class) $\uparrow$ CLIP-T (subject) $\uparrow$
Real Image (on training set) - 0.774 0.885 0.298 0.311
Textual Inversion [7] SD 0.569 0.780 0.255 -
Textual Inversion (our impl) SD 0.611 0.772 0.267 0.289
DreamBooth [25] SD 0.668 0.803 0.305 -
DreamBooth (our impl) SD 0.682 0.808 0.301 0.306
Ours SD 0.704 0.824 0.293 0.309
Ours SDXL 0.744 0.842 0.297 0.312
Table 1: Evaluation on the DreamBench.
Ours (SD) vs DreamBooth Ours (SDXL) vs Ours (SD)
Subject Alignment 64.9% / 35.1% 67.8% / 32.2%
Textual Alignment 52.8% / 47.2% 46.8% / 53.2%
Table 2: Human preference comparison
Figure 3: Rephrasing prompts. Even when using rephrased prompts, our method maintains subject identity, a quality DreamBooth lacks. Notably, SDXL consistently generates images with minimal deviation from the originals.
Figure 4: Influence of adaption to living entities. Without adaptation, the model might overlook motion in the prompts and focus solely on assembling objects within the images.
Figure 5: Overfitting Prevention. Our regularization dataset effectively prevents the model from overfitting to the training images. Interestingly, when we employ a regularization dataset with a simplistic prompt “a [class noun]”, the model exhibits improved text alignment during initial iterations, but experiences a decline in performance over time.
In our quantitative assessment, we follow DreamBooth to employ DINO [4] and CLIP-I [22], to evaluate subject fidelity. For text alignment evaluation, we argue that the CLIP-T used in DreamBooth may yield inaccurate measurements of text-image alignment when vague class names are employed. More details and discussion can be found in Appendix B. To rectify this issue, we replace vague class names with highly specific object names, which results in a substantial improvement in the CLIP-T score for ground truth text-image pairs, as shown in Table 1.
We conducted a user study to compare with DreamBooth [25]. We asked 20 people to answer 30 questions about subject alignment and 30 questions about text alignment, in a total of 1200 questions. The questions are randomly sampled from a large pool of DreamBench test prompts. The details can be found in Appendix C. As shown in Table 2, our method is overwhelmingly preferred for subject alignment and comparable in text alignment. We also compared the effect of a different base model with SDXL instead of SD. SDXL is more expressive in details and hence leads to better subject alignment. For textural alignment SD works slightly better.
4.2 Model Analysis
Does the model overfit to the formatted prompts? To assess the impact of prompt phrasing variations, we rephrase the prompts and compare them to the originals, as illustrated in Figure 3. While our model consistently employs formatted prompts starting with “a <new>[class noun]” during training, it demonstrates resilience against overfitting to this specific format. In contrast, DreamBooth displays a notable departure from subject identity. This observation confirms our assertion that the simplistic use of unvaried prompts in the format “a [class noun]” for constructing a regularization dataset can lead to overfitting to sentence structure. Additionally, it is noteworthy that SDXL exhibits a higher degree of consistency in generating images in response to rephrased prompts, with minimal deviation from the originals.
How important is the formatted prompt? We conducted an ablation test with two variations: 1) without using a regularization dataset, and 2) using a regularization dataset with simple prompts in the format “a [class noun]”. The results are illustrated in Figure 5. These findings emphasize the importance of both the regularization dataset and the utilization of well-structured prompts. Notably, when employing a regularization dataset without well-structured prompts, there is an initial increase in diversity during early training phases compared to scenarios without any regularization dataset. However, over extended training iterations, a noticeable decline in model performance becomes evident. We attribute this decline to the incongruity between the overly simplistic prompt format “a [class noun]” and the complexities of the generated images. This observation underscores the pivotal role of the well-structured prompts in enhancing model performance, and thus highlights their significance in our approach.
Does the adaption to living entities help? The key difference between inanimate objects and living things is motion. Training without adaptation tends to overlook motion while still generating objects separately, as shown in Figure 4. However, it is important to note that adaptation makes generating motion easier, but the model can still produce motion even without it.
How does the training and regularization dataset size affects the performance? We conducted grid experiments on number of training examples 1, 2, 4, and number of regularization examples 100, 500, 1000, 2000. The results are illustrated in Figure  8 in Appendix F. Our method is robust to smaller regularization set size, as only 100 regularization examples can also effectively prevent overfitting and the model is still able to preserve the very fine details of the subjects.
Is the random dropout of adjective words necessary? When altering attributes like color, we found that using random dropout is essential. Without it, the model tends to excessively preserve the original identity, and struggle in color changes. With random dropout, the task becomes easier while still preserving identity, as shown in Figure 9 in Appendix F. It’s worth noting that for color modification, using “a <new>[color] [class noun]” works better than “a [color] <new>[class noun]”.
5 Limitations
Despite its impressive quality, our method has two limitations: 1) It needs to generate a regularization dataset for each category, which increases the overall training time (in practice $<$500 images also work for simple cases); 2) On a single A100 GPU, it requires approximately 1.5 hours for SD and 3.5 hours for SDXL to complete 4000 and 8000 training iterations, respectively. We leave acceleration to future research. Currently, we assume that the real images of the target object are annotated manually. In the future, caption generator could be used, such as BLIP [16, 15].
6 Conclusion
We introduced a data-centric approach to enhance diffusion model personalization. We proposed to incorporate formatted prompts and their generated images to generate a structured regularization dataset. Our method effectively reduces overfitting. This results in significant improvements in image quality, identity preservation, and diversity of generated samples aligned with input text prompts. Although this approach requires extended training times and increased memory resources, we foresee potential solutions in the future, as our method sets the stage for forthcoming data-centric approaches. Notably, our method focuses on data augmentation rather than model architecture adjustments, making it potentially compatible for different pre-trained text-to-image diffusion models. For future work, we would like to focus on 1) generalize to multi-concepts; 2) apply on videos; and 3) accelerate the procedure.
Appendix A Formatted prompt generation
For non-live objects, we use the following prompts to generate words describing shape, color, textures, and background in ChatGPT:
  • shape: give me 100 adjective words describing the shape of an object
  • color: give me 100 adjective words describing the color of an object
  • texture: give me 100 adjective words describing the texture of an object
  • background: give me 500 phrases that describe the background, such as “on the table”, as diverse as possible.
After removing duplicated ones, there are 85 shapes, 93 colors, 96 textures, and 455 backgrounds.
For live objects, we use the following prompts to generate words describing shape, color, textures, and motion in ChatGPT:
  • body: give me 100 adjective words describing the body of an animal
  • skin: give me 100 adjective words describing the skin or fur of an animal
  • emotion: give me 100 adjective words describing the emotion of an animal
  • motion: give me 1000 different short concise sentences that contains a special token “\$concept” which stands for a specific animal, which can be a dog, a cat or a human. For example: “a \$concept sitting in a temple”, “a \$concept walking in a supermarket”. Keep “a \$concept” in the sentences.
After removing duplicated ones, there are 89 bodies, 86 skins/furs, 75 emotions, and 744 motions. For humans, we replace the word ”animal” above with ”person”.
We use
  • style: give me 100 image style descriptions, such as “a photo of”, and “a painting of”.
After removing duplicated ones, there are 99 styles left.
Appendix B Better Category Naming
We show the name change in Table 3. As shown in Figure 6, there is a notable discrepancy between the utilization of vague class names, such as “toy”, and more specific object names, such as “duck toy”, on the ground truth images. Notably, the CLIP-T score appears to be significantly influenced by the nomenclature chosen for the object, and thereby potentially undermines its accuracy as an indicator of text-image alignment. To delve deeper into this matter, we calculate the CLIP-T score on the original images and the manually added prompts. Table 1 presents that, when using vague class names, the CLIP-T score for ground truth text-image pairs falls even below the conventional threshold of 0.3, which is typically considered as the threshold for assessing text-image pair compatibility [28]. To rectify this issue, we replace vague class names with highly specific object names, which results in a substantial improvement in the CLIP-T score for ground truth text-image pairs. Additional details regarding the modified nomenclature list can be found in Appendix B.
Figure 6: CLIP-T score on different subject names. We randomly generate an image with the prompt “a teddy bear toy sitting beside a river”. Subsequently, we evaluate prompts “a {} in the grass next to a tree” and “a {} sitting on a sofa”. When employing the vague prompt “toy”, both the ground truth and the mismatched example are classified as mismatches (<0.3). In contrast, when using specific class names, the ground truth is categorized as a match (>0.3), while the mismatched examples remain classified as mismatches.
subject name original class modified class
bear_plushie stuffed animal bear plushie
berry_bowl bowl berry bowl
can can drink can
clock clock alarm clock
duck_toy toy duck toy
grey_sloth_plushie stuffed animal sloth plushie
monster_toy toy monster toy
poop_emoji toy poop emoji toy
rc_car toy racing car toy
red_cartoon cartoon 2d cartoon devil
robot_toy toy robot toy
wolf_plushie stuffed animal wolf plushie
Table 3: Name Change. We change the name for a more reasonable CLIP-T metric and better performance.
Appendix C Details of User Study
We randomly sampled and paired 300 comparisons of ours(SD) versus DreamBooth, half of which is for the subject alignment and the other half for the text alignment. For subject alignment, we randomly sampled a ground truth image and asked ”The foreground object in which image is more similar to the reference?”. For text alignment, we asked ”Which image better depicts {}?”, where {} is replaced by the prompt. We equally divided the questions into 10 groups. Each person randomly received one group. We did the same for ours(SDXL) versus ours(SD). We provide an example of our interface in Figure 7.
Figure 7: An example of the user study interface. The left is an example for question of object alignment, and the right is an example for question of text alignment. Each user was asked to answer 30 questions about subject alignment (left) and 30 questions about text alignment (right). The examples in the questions are randomly sampled from a large pool.
Appendix D Extension: Using BLIP to Generate Captions
We tried to use BLIP [16] to generate more personalized captions for the training example images. BLIP outputs a caption for the input image and can be conditioned on the format. We condition BLIP so that it generates prompts that start with ”a [subject]”. For instance for the ”tortoise plushie” image BLIP generates
  • a tortoise plushie on a pillow
  • a tortoise plushie
  • a tortoise plushie sitting on a piano keyboard
  • a tortoise plushie on a desk
To unify the prompt format, we task ChatGPT to‘Change the following sentence to the format ”A $<$new$>$ tortoise plushie blablabla”. The ”$<$new$>$” is a special token that needs to be inserted before tortoise plushie.’ The result are the following prompts
  • a $<$new$>$ tortoise plushie on a pillow
  • a $<$new$>$ tortoise plushie
  • a $<$new$>$ tortoise plushie sitting on a piano keyboard
  • a $<$new$>$ tortoise plushie on a desk
The results of this ”tortoise plushie” dataset is shown in Figure 15. With this addition of using BLIP, it alleviated writing the prompt examples manually, i.e., it replaced the manual steps in Section 3.
Appendix E Implementation Details
We opt for the identifier word “olis” instead of the more commonly used “sks”. This choice is based on the fact that “olis” corresponds to the least frequently utilized token in the model’s vocabulary [1]. Each training batch contains one example from training set and one example from regularization set. For SD, we fine-tune the entire model with a learning rate of 2e-6 and perform inference using 200 steps of DDIM [31]. For SDXL, which has a larger model size, we employ a LoRA with a rank of 32 for both the text encoders and UNet. We also train the text embeddings. We set learning rate to 1e-4. We use 50 steps of DDIM for inference. We show the best number of iterations in Table 4. For simplicity, we use 4000 and 8000 iterations for SD and SDXL, respectively.
subject name best #iterations on SD best #iterations on SDXL
backpack 6000-8000 8000-10000
backpack_dog 2000-3000 4000-6000
bear_plushie 2000-4000 4000-6000
berry_bowl 6000-8000 8000-10000
can 6000-8000 8000-10000
candle 4000-6000 8000-10000
cat 1000-3000 1000-3000
cat2 6000-8000 8000-10000
clock 6000-8000 8000-10000
colorful_sneaker 4000-6000 6000-8000
dog 1000-3000 1000-3000
dog2 2000-4000 4000-6000
dog3 2000-4000 8000-10000
dog5 3000-4000 6000-8000
dog6 3000-4000 6000-8000
dog7 3000-4000 6000-8000
dog8 1000-3000 1000-3000
duck_toy 3000-4000 3000-4000
fancy_boot 3000-4000 6000-8000
grey_sloth_plushie 3000-4000 6000-8000
monster_toy 3000-4000 8000-10000
pink_sunglasses 3000-4000 4000-6000
poop_emoji 3000-4000 4000-6000
rc_car 3000-4000 4000-6000
red_cartoon 6000-8000 8000-10000
robot_toy 3000-4000 6000-8000
shiny_sneaker 3000-4000 6000-8000
teapot 6000-8000 8000-10000
vase 6000-8000 8000-10000
wolf_plushie 3000-4000 4000-6000
Table 4: Best #iterations of datasets in DreamBench. The variation mainly comes from the diversity of the dataset itself.
Appendix F More Results
Figure 8: Ablation Tests on number of training examples and regularization dataset size. Even only a very small regularization dataset is given (1̃00 examples), our method still effectively prevent overfitting and preserves the identity.
Figure 9: Color Modification. Our method can alter the color of the subject. It is important to mention that when modifying the color, using “a <new>[color] [class noun]” is more effective than “a [color] <new>[class noun]”.
Figure 10: More Results on inanimate objects.
Figure 11: More Results on inanimate objects.
Figure 12: More Results on living entities.
Figure 13: More Results on living entities.
Figure 14: More Comparison.
Figure 15: More Comparison.
Figure 16: More Comparison.